# CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes

### v4

@@ 3.3 Results and comparisons
CD-NGP 21M 30.23 0.198 75min 113MB
3D+4D 23M 29.34 0.243 82min 122MB
CD-NGP-C 21M 30.52 0.174 80min 114MB
CD-Plane 14M 29.16 0.215 77min 83MB
CD-MERF 7.5M 29.93 0.201 109min 57MB
CD-MERF-C 7.5M 30.22 0.174 115min 57MB
Our proposed continual learning framework is compatible with multiple spatial backbones. We provide quantitative results on hybrid representation (denoted as CD-MERF) and plane representation like [17, 21] (denoted as CD-Plane) on the DyNeRF dataset in Table 3. Since the element-wise product for plane representation in [17] produces unsatisfactory results due to the hash collision, we remove the $P_{m}(x,y,z)$ in Eq. 6 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation. We use $2^{P-2}$-length hash tables for the Plane representation in Table 3 to balance the quality and the model size. The structural configurations $(L,F)$ of the plane representation in Table 3 are set as $(6,4)$ for better results. The 3D+4D representation similar to [34] is yielded by concatenating the spatial $(x,y,z)$ features and spatial-temporal $(x,y,z,t)$ features encoded by a 4D extension of Eq. 5. By comparing our CD-NGP with the naive 3D+4D representation, we show the advantage of our asymmetric encoding of spatial-temporal $(x,y,z,t)$ features explained in Eqs. 3 and 7: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and lead to better results.
Table 3 also shows metrics of different feature fusion methods. We replace the element-wise sum in Eqs. 7 and 6 with vector concatenation and denote them as CD-NGP-C and CD-MERF-C in Table 3. The hybrid representations (CD-MERF and CD-MERF-C) achieve comparable model sizes ($\sim$60MB) to methods for static scenarios [31, 21] without significant quality loss, reducing the average bandwidth cost to $0.2$MB/frame. Benefiting from larger feature dimensions, the vector concatenation leads to slightly better quality (especially when measured in LPIPS) by more training time, and the reasonable trade-off also shows the robustness of our method. Compared with the hash table designs in Table 7, the CD-NGP-C representation in Table 3 almost performs as well as the bulky $(18,18)$ configuration in Table 7 and reduces the model size by $80\%$. We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Table 4.
Model Flame Salmon Cook Spinach
[…]
CD-NGP 26.37 0.04586 0.2886 30.76 0.02440 0.1793
CD-NGP-C 26.65 0.04243 0.2680 31.03 0.02288 0.1593
CD-Plane 25.48 0.06058 0.3173 29.41 0.03487 0.2010
CD-MERF 26.15 0.04952 0.2921 30.35 0.02931 0.1858
CD-MERF-C 26.56 0.04435 0.2546 30.69 0.02609 0.1583
Model Flame Steak Sear Steak
[…]
CD-NGP 31.27 0.02132 0.1736 31.71 0.01917 0.1701
CD-NGP-C 31.64 0.01823 0.1403 32.02 0.01698 0.1424
CD-Plane 30.26 0.02949 0.1804 30.56 0.02723 0.1811
CD-MERF 31.01 0.02508 0.1719 31.44 0.02181 0.1698
CD-MERF-C 31.13 0.02424 0.1506 31.66 0.01989 0.1413
Model Cut Roasted Beef Average
[…]
CD-NGP 31.03 0.02366 0.1801 30.228 0.026882 0.19834
CD-NGP-C 31.26 0.02205 0.1578 30.520 0.024514 0.17356
CD-Plane 30.11 0.03302 0.1951 29.164 0.037038 0.21498
CD-MERF 30.71 0.02765 0.1856 29.932 0.030674 0.20104
CD-MERF-C 31.04 0.02450 0.1632 30.216 0.027814 0.17360
@@
The detailed parameter numbers and bandwidth consumption are listed in Table 5. 3D Enc. and 2D Enc. denote the spatial encodings along voxels and planes respectively. The tuples $(M_{1},M_{2})$ in the cells denote the parameter numbers of the base encoder and each of the auxiliary encoders. Following [27, 43, 1, 48] we also report the minimum per-frame bandwidth consumption $\mathbf{B_{min}}$, which only depends on the parameter numbers of the auxiliary branches and is thus significantly lower than the average bandwidth consumption $\mathbf{B_{avg}}$. Since the base encoder and the density grid could be shared among different branches and transmitted only once in the beginning, the minimum bandwidth consumption is reduced to $0.065\sim 0.16$ MB per frame, which is 100×100$\times$100 × less than the voxel-based StreamRF [27] and 7×7$\times$7 × less than the pure MLP representation INV [48].
Table 5: The parameter numbers of each module of the 30-branch representations in Table 3. $\mathbf{B_{min}}$ and $\mathbf{B_{avg}}$ are minimum and average bandwidth costs in MB/frame. ##\## denotes the number of parameters per branch (10 frames).
Model ##\##3D Enc. ##\##2D Enc. ##\##MLP ##\##Online ##\##Total $\mathbf{B_{min}}$ $\mathbf{B_{avg}}$
CD-NGP $(9.9M,0.37M)$ $0.042M$ $0.39M$ $21M$ 0.16 0.38
CD-NGP-C $(9.9M,0.37M)$ $0.046M$ $0.40M$ $21M$ 0.16 0.38
CD-Plane - $(1.9M,0.082M)$ $0.048M$ $0.28M$ $14M$ 0.11 0.28
CD-MERF $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.042M$ $0.16M$ $7.5M$ 0.061 0.19
CD-MERF-C $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.052M$ $0.17M$ $7.5M$ 0.065 0.19
@@
We also provide quantitative comparisons on our proposed long multi-view video dataset in Table 6 and Fig. 7 to further show the superior scalability of our method. We split the sequence into 300-frame chunks for the compared methods since the offline baseline [7] is designed for 300-frame sequences in the DyNeRF dataset. We set the hash table size as $2^{12}$ in the auxiliary branches and use $T_{chunk}=5$ in our model. As shown in Table 6, our method reduces the memory and storage consumption by at least 4×4$\times$4 × with comparable quality to HexPlane [7]. Compared with the results in Table 1, the performance gap in storage consumption is even larger in long video sequences. This is because the whole model shares one base branch and the following auxiliary branches occupy the majority of the parameters. We thus set the hash table size in the auxiliary branches to $2^{12}LF$ and leverage the re-use of the features from the base branch to reduce the model size to $\mathcal{O}(N_{chunk}\cdot 2^{12}LF)$. .
The PSNR curves in Fig. 7 further demonstrate the stability of our method. The compared online methods suffer from catastrophic forgetting: the PSNRs at the beginning of the chunks are much lower than that in the end. The problem is properly eliminated by our method through parameter isolation, and our method continuously produce high-quality results for the whole sequence. We also provide comparisons between the rendered results in Fig. 8. The regularization method displayed in Fig. 8(a) reconstructs the TV screen in the background erroneously. The replay method displayed in Fig. 8(b) fails to recover the details. Ours in Fig. 8(f) produces comparable quality to the offline baseline in Fig. 8(c).
To further demonstrate the scalability of our method, we conduct experiments on the first 2400 frames of the seminar scene in our long video dataset, show the PSNR curves in Fig. 10 and compare the model sizes in Fig. 11. For simplicity of the evaluation, we set the chunk size as 10 for the CD-NGP and CD-MERF method. Our method produces high-quality results, and consumes only a fraction of storage space compared with the other online baselines.
Figure 7: PSNR curves of the first 1200 frames of seminar scene in our long video dataset. The solid lines and dotted lines denote offline and online methods respectively.
@@ 3.4 Ablation studies
Figure 9: PSNR curves of asymmetric (solid line, ours) and symmetric (dotted line) hash table designs on the sear steak scene.
Figure 10: PSNR curves of the first 2400 frames of seminar scene in our long video dataset.
Figure 11: Comparison of the model sizes of our method and other online baselines. Our method show superior scalability over the compared baselines.
@@
Benefiting from the multi-branch hash table design, our model is highly efficient and configurable. We change the parameter numbers of the initial branch and the auxiliary branches and obtain the results on the DyNeRF dataset in Table 7. $(P_{1},P_{2})$ are the log2 values of the hash table lengths of the base branch and the other branches. Since the whole model contains hash encoders, MLP regressors, and the occupancy grids. The sizes of the models are thus larger than the simple combinations of hash table and MLP regressors. We use the default configuration $(19,14)$ in the following experiments. Although larger hash tables alleviate hash collisions, the configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40\%$ lower. Moreover, the $(19,12)$ configuration of CD-NGP significantly outperforms the naive $(14,14)$ configuration at a similar storage cost. To further demonstrate the scalability of our method, we provide the PSNR curves of different hash table designs on the sear steak scene in Fig. 9. Equipped with a larger hash table in the base branch, the asymmetric setting $(19,14)$ and $(19,12)$ outperforms others in the beginning and produces comparable metrics to the naive $(16,16)$, $(14,14)$ configuration respectively in the subsequent chunks. The advantage comes from our asymmetric hash table designs: the base branch with a large hash table recovers more environment patterns, and the auxiliary branches could thus use fewer parameters to cope with recent changes only.
@@
HexPlane $\checkmark$ 28.32 0.117 800MB 8 hours 74GB
HexPlane-PR $\checkmark$ 20.90 0.286 1272MB 8 hours 12GB
HexPlane-GR $\checkmark$ 24.82 0.282 1272MB 8 hours 12GB
StreamRF $\checkmark$ 24.99 0.266 13GB 4 hours 12GB
INV $\checkmark$ 27.07 0.086 1344MB 160 hours 12GB
CD-NGP (Ours) $\checkmark$ 27.15 0.190 201MB 10 hours 12GB
@@
We also conduct ablation studies on the field composition methods and show the results in Tables 9 and 12. The field composition process is defined as:
@@
Full composition in Table 9 denotes the naive composition of 2 fields from the base branch and the current branch respectively, i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ are the field from the base branch and $\sigma_{a}$ and $\mathbf{c}_{a}$ in Eq. 13 are the field from the current branch. Both branches are encoded with the same architecture as Eq. 3. Static-dynamic in Table 9 denotes a naive composition of the field from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [45], i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ in Eq. 13 are encoded without the $\Gamma(t)$ in Eq. 3. The online training strategies remain the same. Our CD-NGP outperforms the other methods in all metrics. Furthermore, as compared in Fig. 12, the compared composition methods in Figs. 12(c) and 12(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.
@@
3.4.4 Regularization weights.
We also provide ablation studies on the regularization weights of auxiliary spatial encodings. We apply regularizations on $\Psi_{k}(x)$ with different $\lambda_{r}$ and show the results in Table 10. The regularization is of slight improvement, and our method is robust within a wide range of $\lambda_{r}$.
Table 10: Ablation study on regularization weight $\lambda_{r}$.
[…]
$10^{-5}$ 30.20 0.0269 0.198
$10^{-4}$ 30.18 0.0268 0.196
$10^{-3}$ 30.23 0.0269 0.198
$10^{-2}$ 30.09 0.0279 0.197
1 29.66 0.0313 0.205
3.4.5 Chunk size.
We also conduct ablation studies on the chunk sizes and show the results in Tables 11 and 12. We only report the results on the cook spinach scene in the DyNeRF dataset for brevity. Our method is robust within a wide range of chunk sizes. As in Table 11, the smaller chunk size leads to better results when controlling the iteration numbers of each chunk. When the total iterations of the whole 300-frame sequence are held constant, the larger chunk size leads to better results as in Table 12. We select $T_{chunk}=10$ as the default configuration. The selection of the chunk size is a trade-off between the quality, robustness, training time, and host memory occupation and should depend on the data.
Table 11: Ablation study on chunk sizes. The iteration numbers of each chunk are held constant under different $T_{chunk}$ settings.
@@
Table 12: Ablation study on chunk sizes. The total iterations of the whole 300-frame sequence are held constant under different $T_{chunk}$ settings.
@@ 4 Conclusion and Limitation
We propose CD-NGP, a continual learning framework for multi-view dynamic scenes based on parameter isolation. Equipped with the multi-branch hash encoding structure, it shows high efficiency and scalability and is capable of reconstructing long video sequences with low host memory occupation. Compared with offline methods, it consumes $40\%$ less bandwidth and $85\%$ less memory than recent explicit baselines and achieves comparable quality while maintaining the speed advantage over the pure-MLP-based methods. Compared with online baselines, it significantly outperforms replay and regularization-based continual learning methods and reaches a new balance among reconstruction quality, training speed, bandwidth occupation, and memory cost. However, like most of the NeRF-based methods, it still requires querying the field function and is not appropriate for achieving real-time rendering. Further efforts could be made to apply baking methods or Gaussian point clouds to continual learning to achieve real-time rendering in long video sequences for wide applications.

### v5

@@ 3.3 Results and scalability analysis
CD-NGP 21M 30.23 0.198 75min 113MB $(9.9M,0.37M)$ $0.39M$ 0.16 0.38
3D+4D 23M 29.31 0.242 82min 122MB $(0.73M,0.73M)$ $0.76M$ 0.31 0.41
CD-NGP-C 21M 30.52 0.174 80min 114MB $(9.9M,0.37M)$ $0.40M$ 0.16 0.38
CD-Plane 14M 29.16 0.215 77min 83MB - $(1.9M,0.082M)$ $0.28M$ 0.11 0.28
CD-MERF 7.5M 29.93 0.201 109min 57MB $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.16M$ 0.061 0.19
CD-MERF-C 7.5M 30.22 0.174 115min 57MB $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.17M$ 0.065 0.19
@@
The parameter numbers and bandwidth consumption in Table 3 also prove the superior scalability of our method, where 3D Enc. and 2D Enc. denote the spatial encodings according to voxel and plane indices respectively. Following MERF [38] for static scenes, we use a smaller hash table for the hybrid representation in Eq. 12 than pure voxel hash grid [31] representation in Table 3 to ensure the compactness and scalability of our method. Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the CD-MERF are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively. The number of the parameters of the MERF representation is reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids. Since the hash tables comprise most of the parameters, the model size is reduced from $21M$ to $7.5M$ as listed in Table 3. Likewise, we use $2^{P-2}$-length hash tables for the Plane representation in Table 3 to balance the quality and the model size. The structural configurations $(L,F)$ of the plane representation in Table 3 are set as $(6,4)$ , and the resolution of the hash encoders of the CD-Plane are set within $(64,2048)$ for better results. Following [27, 43, 1, 48] we also report the minimum per-frame bandwidth cost $\mathbf{B_{min}}$, which only depends on the parameter numbers of the auxiliary branches and is thus significantly lower than the average bandwidth consumption $\mathbf{B_{avg}}$. Since the base encoder and the density grid could be shared among different branches and transmitted only once in the beginning, the minimum bandwidth consumption is reduced to $0.065\sim 0.16$ MB per frame, which is 100×100$\times$100 × less than the voxel-based StreamRF [27] and 7×7$\times$7 × less than the pure MLP representation INV [48].
@@
As introduced in Section 2.3, our method shows superior scalability: the total size of the model is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, where $N_{chunk}$ is the number of chunks in the sequence. To prove the scalability of our method, we provide quantitative comparisons on our proposed long multi-view video dataset in Table 4 and Fig. 6(b) to further show the superior scalability of our method. We split the sequence into 300-frame chunks for the compared methods since the offline baseline [7] is designed for 300-frame sequences in the DyNeRF dataset. We set the hash table size as $2^{12}$ in the auxiliary branches and use $T_{chunk}=5$ in our model. As shown in Table 4, our method reduces the memory and storage consumption by at least 4×4$\times$4 × with comparable quality to HexPlane [7]. Compared with the results in Table 1, the performance gap in storage consumption is even larger in long video sequences. This is because the majority of the parameters in our method are used for the auxiliary branches, and the base branch only takes up a small fraction of the parameters. Consequently, applying smaller hash tables in the auxiliary branches and leverage the re-use of the features in the base branch leads to a significant reduction in the model size.
The PSNR curves in Fig. 6(b) further demonstrate the stability of our method. The compared regularization-based and replay-based methods suffer from catastrophic forgetting: the PSNRs at the beginning of the chunks are much lower than those in the end. The problem is properly eliminated by our method through parameter isolation, and our method continuously produces high-quality results for the whole sequence. We also provide comparisons between the rendered results in Fig. 7. The regularization method displayed in Fig. 7(a) reconstructs the TV screen in the background erroneously. The replay method displayed in Fig. 7(b) fails to recover the details. Ours in Fig. 7(f) produces comparable quality to the offline baseline in Fig. 7(c).
To further demonstrate the scalability of our method, we conduct experiments on the first 2400 frames of the seminar scene in our long video dataset. We further show the PSNR curves in Fig. 6(c) and compare the model sizes in Fig. 8. Our method produces high-quality results while consuming only a fraction of storage space compared with the other online baselines.
@@ 3.4 Ablation studies
Figure 8: Comparison of the model sizes of our method and other online baselines. Our method shows superior scalability over others.
@@
Benefiting from the multi-branch hash table design, our model is highly efficient and configurable. We change the lengths of the hash tables of the initial branch and the auxiliary branches and obtain the results on the DyNeRF dataset in Table 5. $(P_{1},P_{2})$ are the log2 values of the hash table lengths of the base branch and the auxiliary branches. The whole model contains hash encoders, MLP regressors, and the occupancy gridd, the sizes of the models are thus larger than the value deduced from the combinations of the hash table and MLP regressors. Although larger hash tables lead to less hash collisions, the default configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40\%$ lower. Moreover, the $(19,12)$ configuration of CD-NGP significantly outperforms the naive $(14,14)$ configuration at a similar storage cost. To further demonstrate the superiority of the asymmetric hash table design and prove the scalability of our method, we provide the PSNR curves of different hash table designs on the sear steak scene in Fig. 6(d). Equipped with a larger hash table in the base branch, the asymmetric setting $(19,14)$ and $(19,12)$ outperforms others in the beginning and produces comparable metrics to the naive $(16,16)$, $(14,14)$ configuration respectively in the subsequent chunks. The $(18,18)$ in Table 5 configuration delivers the best results (30.55 dB in PSNR) on the DyNeRF dataset, but at the cost of an exceptionally large model size (638MB). However, the CD-NGP-C configuration with $(P_{1},P_{2})=(19,14)$ listed in Table 3 achieves a comparable PSNR (30.52 dB) with a much smaller model size (114MB, $80\%$ compression). The advantage comes from our re-use of the spatial features encoded by different branches: the base branch with a large hash table recovers more environment details, and the auxiliary branches could thus use fewer parameters to cope with recent changes only. The combination of the base encode and the auxiliary encoders lead to a more efficient and scalable model.
@@
MixVoxels $\checkmark$ 26.94 0.022 0.118 2000MB 80 min 80GB
HexPlane $\checkmark$ 27.74 0.022 0.117 800MB 8 hours 74GB
HexPlane-PR $\checkmark$ 20.90 0.095 0.286 1272MB 8 hours 12GB
HexPlane-GR $\checkmark$ 24.82 0.052 0.282 1272MB 8 hours 12GB
StreamRF $\checkmark$ 24.72 0.036 0.273 13GB 4 hours 12GB
INV $\checkmark$ 27.12 0.025 0.086 1344MB 160 hours 12GB
CD-NGP (ours) $\checkmark$ 27.15 0.039 0.188 201MB 10 hours 12GB
@@
We also conduct ablation studies on the field composition methods and show the results in Tables 7 and 9. The field composition process is defined as:
@@
Full composition in Table 7 denotes the naive composition of 2 fields from the base branch and the current branch respectively, i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ are the field from the base branch and $\sigma_{a}$ and $\mathbf{c}_{a}$ in Eq. 14 are the field from the current branch. Both branches are encoded with the same architecture as Eq. 3. Static-dynamic in Table 7 denotes a naive composition of the field from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [45], i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ in Eq. 14 are encoded without the $\Gamma(t)$ in Eq. 3. The online training strategies remain the same. Our CD-NGP outperforms the other methods in all metrics. Furthermore, as compared in Fig. 9, the compared composition methods in Figs. 9(c) and 9(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.
@@
Table 8: Ablation study on the chunk sizes. The iteration numbers of each chunk are held constant under different $T_{chunk}$ settings.
@@
Table 9: Ablation study on the chunk sizes. The total iterations of the whole 300-frame video are held constant under different $T_{chunk}$ settings.
@@
3.4.4 Chunk size.
We also conduct ablation studies on the chunk sizes and show the results in Tables 8 and 9. We only report the results on the cook spinach scene in the DyNeRF dataset for brevity. As in Table 8, the smaller chunk size leads to better results when controlling the iteration numbers of each chunk. When the total iterations of the whole 300-frame sequence are held constant, the larger chunk size leads to better results as in Table 9. The selection of the chunk size is a trade-off between the quality, robustness, training time, and host memory occupation and should depend on the data, and our method is robust within a wide range of chunk sizes. We select $T_{chunk}=10$ as the default configuration.
@@ 4 Conclusion and Limitation
We propose CD-NGP, a continual learning framework for multi-view dynamic scenes based on parameter isolation, along with a self-collected dataset for novel view synthesis under multi-view long-view scenarios. Our proposed method splits the input multi-view videos into different chunks, trains different model branches, and performs reconstruction accordingly. Each branch encodes spatial and temporal features separately using the multi-resolution hash encoding structure. It leverages the redundancy of the dynamic scenes and reuses the spatial features in the first branch to achieve fast convergence, high fidelity, and superior scalability. On the prevailing DyNeRF dataset, it consumes $40\%$ less bandwidth and $85\%$ less memory than recent explicit offline baselines, achieves comparable quality, and maintains the speed advantage over pure-MLP-based offline methods. Compared with online baselines, it delivers state-of-the-art quality and achieves a new balance among reconstruction quality, training speed, bandwidth occupation, and memory cost. On the long video dataset, our method delivers high quality and achieves fast convergence with only a fraction of the storage cost. However, like most NeRF-based methods, it still requires querying the field function and is not suitable for real-time rendering. Further efforts could be made to apply baking methods or Gaussian point clouds to continual learning, enabling real-time rendering in long video sequences for wide applications.
Appendix A Per-scene results
We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Table 10, and the per-scene results of the methods in Table 4 on our self-collected long video dataset in Table 12.
Model Flame Salmon Cook Spinach Cut Roasted Beef
[…]
CD-NGP 26.37 0.046 0.289 30.76 0.024 0.179 31.03 0.024 0.180
3D+4D 25.72 0.059 0.355 29.73 0.032 0.220 30.06 0.031 0.224
CD-NGP-C 26.65 0.042 0.268 31.03 0.022 0.159 31.26 0.022 0.158
CD-Plane 25.48 0.060 0.317 29.41 0.034 0.201 30.11 0.033 0.195
CD-MERF 26.15 0.049 0.292 30.35 0.029 0.186 30.71 0.028 0.186
CD-MERF-C 26.56 0.044 0.255 30.69 0.026 0.158 31.04 0.025 0.163
Model Flame Steak Sear Steak Average
[…]
CD-NGP 31.27 0.021 0.174 31.71 0.019 0.170 30.23 0.027 0.198
3D+4D 30.33 0.028 0.205 30.73 0.026 0.207 29.31 0.035 0.242
CD-NGP-C 31.64 0.018 0.140 32.02 0.017 0.142 30.52 0.025 0.174
CD-Plane 30.26 0.029 0.180 30.56 0.027 0.181 29.16 0.037 0.215
CD-MERF 31.01 0.025 0.172 31.44 0.022 0.170 29.93 0.031 0.201
CD-MERF-C 31.13 0.024 0.151 31.66 0.020 0.141 30.22 0.028 0.174
Table 11: Per-scene results of the compared methods on the meetroom dataset in Table 2.
Model Trimming Discussion
[…]
MixVoxels 27.21 0.026 0.093 27.42 0.026 0.110
HexPlane 22.13 0.062 0.263 26.28 0.032 0.212
HexPlane-PR 17.30 0.140 0.404 18.53 0.111 0.355
HexPlane-GR 12.80 0.184 0.554 23.31 0.061 0.330
StreamRF 26.72 0.029 0.170 26.65 0.028 0.169
INV 27.08 0.026 0.089 24.57 0.036 0.110
CD-NGP 27.58 0.040 0.164 26.01 0.040 0.214
Model VRheadset Average
[…]
MixVoxels 26.25 0.029 0.107 26.96 0.027 0.103
HexPlane 25.72 0.035 0.187 24.71 0.043 0.221
HexPlane-PR 20.05 0.085 0.317 18.63 0.112 0.359
HexPlane-GR 22.37 0.054 0.316 19.49 0.099 0.400
StreamRF 26.06 0.029 0.171 26.48 0.029 0.170
INV 19.95 0.061 0.160 23.87 0.041 0.120
CD-NGP 24.43 0.037 0.193 26.01 0.039 0.191
Table 12: Detailed results of the proposed CD-NGP and other hash-based methods on our long video dataset.
Model Demo Discussion Kungfu
[…]
MixVoxels 27.50 0.021 0.103 28.01 0.016 0.097 24.43 0.033 0.140
HexPlane 28.92 0.016 0.095 27.17 0.024 0.136 25.14 0.032 0.136
HexPlane-PR 21.10 0.090 0.268 20.98 0.094 0.315 18.85 0.112 0.309
HexPlane-GR 24.85 0.041 0.243 24.03 0.050 0.296 22.15 0.072 0.336
StreamRF 24.86 0.032 0.233 24.41 0.029 0.258 23.75 0.043 0.301
INV 26.55 0.025 0.093 28.51 0.017 0.054 24.57 0.037 0.120
CD-NGP 26.66 0.030 0.161 29.66 0.057 0.137 23.66 0.055 0.257
Model Reading Seminar Average
[…]
MixVoxels 25.99 0.021 0.144 28.78 0.018 0.108 26.94 0.022 0.118
HexPlane 27.88 0.018 0.099 29.59 0.018 0.120 27.74 0.022 0.117
HexPlane-PR 21.85 0.087 0.287 21.78 0.092 0.280 20.91 0.095 0.292
HexPlane-GR 26.52 0.040 0.266 25.77 0.053 0.306 24.66 0.051 0.289
StreamRF 24.06 0.040 0.325 26.53 0.037 0.250 24.72 0.036 0.273
INV 26.39 0.026 0.106 29.60 0.018 0.054 27.12 0.025 0.085
CD-NGP 26.70 0.031 0.212 28.97 0.024 0.169 27.13 0.039 0.201